OpenStudy (anonymous):

When you have a trig function like tan^3[x(y^2)+y] and you need to take the derivative, do you still bring down the exponent on the tan function (3tan^2) before you continue?

OpenStudy (noelgreco):

Yes. It's actually easier to see if you use a form like this: \[[\tan(xy ^{2}+y)]^{3}\]

OpenStudy (anonymous):

Afterward would you leave it as \[3[\tan(xy^2+y)]^2\] and keep going?

OpenStudy (noelgreco):

Yep - chain rule, implicit differentiation, and you're there.

OpenStudy (anonymous):

Now when doing the implicit differentiation, do you write: \[d/dx[3\tan(xy^2+y)^2]=dx/dx (x)\] or \[dy/dx[3\tan(xy^2+y)^2]=dx/dx(x)\]

OpenStudy (noelgreco):

\[3[\tan(xy ^{2}+y]^{2}\frac{ dy }{ dx }\tan(xy ^{2}+y)\]

OpenStudy (noelgreco):

That should be d/dx not dy/dx

OpenStudy (anonymous):

Ok cool, it kept throwing me off. Thanks!

OpenStudy (anonymous):

Ok Im working through it and as Im differentiating, when I factor out the dy/dx, that d/dx gets pulled out also?

OpenStudy (noelgreco):

No, that just indicates you must take the derivative of the statement after. The next step would be\[\sec ^{2}(xy ^{2}+y)\] times the derivative of \[(xy ^{2} +y)\]